OpenStudy (anonymous):

convert y= 51.74e^(-0.8x) into a logarithmic function.

OpenStudy (anonymous):

y/51.74=e^(-0.8x) ln(y/51.74)=-0.8x x=(ln(y/51.74))/-0.8

OpenStudy (anonymous):

i have seen these question before what exactly does it mean to turn an exponential function in to a logarithmic function?? it makes no sense whatsoever

OpenStudy (anonymous):

it looks confusing, but it's correct. I'm taking a class on it.

OpenStudy (anonymous):

if the question is "find the inverse function" then that makes sense

OpenStudy (anonymous):

but you're not finding the inverse, you're finding the logarithmic.

OpenStudy (anonymous):

that is like saying "turn this quadratic function in to a cubic function" or "turn this tangent function in to a rational function" nonsense

OpenStudy (anonymous):

they teach nonsense in schools now.

OpenStudy (anonymous):

you found the inverse function if that is what the question is, "find the inverse" then it makes sense otherwise not

OpenStudy (anonymous):

that's actually another way of saying it... it just means to put a log or ln in the equation somehow.

OpenStudy (anonymous):

no it is not another way of saying it, any more than "convert \(\sin(x)\) in to \(\sin^{-1}(x)\)" is the same as saying "find the inverse"

OpenStudy (anonymous):

so is x=(in(y/51.74)/ -0.8 in log form or no?